OpenStudy (surana):

A laboratory employee is mixing a 10% saline solution with a 4% saline solution. How much of each solution is needed to make 500 milliliters of a 6% solution

OpenStudy (surana):

Possible Answers: A: 170 milliliters of the 10% solution and 330 milliliters of the 4% solution. B: 167 milliliters of the 10% solution and 333 milliliters of the 4% solution. C: 110 milliliters of the 10% solution and 390 milliliters of the 4% solution. D: 155 milliliters of the 10% solution and 345 milliliters of the 4% solution.

OpenStudy (hossamhoussien):

Let "x" represent the number of mL of 10% saline solution required to be added. Let "y" represent the number of mL of 4% saline solution required to be added. \[x+y=500\] Now multiply each volume by its concentration \[(\frac{ 10 }{ 100 } \times x)+(\frac{ 4 }{ 100 } \times y)=(\frac{ 6 }{ 100 } \times 500)\] Solving this for x and y you get that: \[x=166.66666 \approx 167ml\] \[y=333.33333 \approx 333ml\]

OpenStudy (mathmale):

@HossamHoussien : It's obvious that you know your subject matter. However, the goal here is to engage your student by guiding that person through the solution of the problem he or she posted. Please do NOT do problems for others and give away the answers. Please read: http://openstudy.com/code-of-conduct Thank you for your interest and involvement in OpenStudy.

OpenStudy (hossamhoussien):

I got it but actually i'm a little bit confused I did post the answer WITH explanation not just the final answer and i meant to explain how it goes with example, so I didn't meant any harm

OpenStudy (amistre64):

If someone does not understand how to get from point A to point B, and a person comes along and does all the driving for them .... the person will always get lost when they try to do it alone. A detailed explanation is of no practical use. If you sit in the passenger seat, you can then have a conversation with the person and let them do the driving ... point out special landmarks or trains of thought along the way that encourage the person to remember the route to take when they are by themself. The key is, THEY have to make it from point A to point B - they have to participate in a meaningful manner.

OpenStudy (surana):

Uh...thanks? It was a detailed explanation, I must admit. I"ll just close this now, since I solved it earlier before I remembered to look at this question to see if it had been solved (and got it wrong because I didn't think to check here to get a second opinion).